Please i need help in solving this quadratic equation 3(5c - 1) = 4(3c + 2)
Answer by 1998Baby(63)

3(5c - 1) = 4(3c + 2)
first distribute
15c-3=12c+8
than add 3 to 8
15c=12c+11
then subtract 12c to 15c
3c=11
than divide 3 to 11
c=11/3